OpenStudy (rayford):

What are the solutions of 5x2 + 2x – 8 = 0?

OpenStudy (helder_edwin):

when u have a cuadratic equaion \(ax^2+bx+c=0\) the "easiest" way to solve it is using the cuadratic formula \[ \large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \] can u do it?

OpenStudy (rayford):

first of all your picture is amazing.. and second of all no that's why i asked for help :[

OpenStudy (helder_edwin):

in your equation u have a=5 b=2 and c=-8 then replacing in the formula u get \[ \large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}= \frac{-2\pm\sqrt{5^2-4\cdot5\cdot(-8)}}{2\cdot5} \]

OpenStudy (helder_edwin):

ok so far?

OpenStudy (rayford):

yes im with you

OpenStudy (helder_edwin):

great

OpenStudy (helder_edwin):

now \[ \large =\frac{-2\pm\sqrt{25+160}}{10}=\frac{-2\pm\sqrt{185}}{10} \]

OpenStudy (rayford):

ok, thank god i'm understanding...

OpenStudy (helder_edwin):

this is it

OpenStudy (helder_edwin):

u have two solutions \[ \large x_1=\frac{-2+\sqrt{185}}{10}\qquad\text{and}\qquad x_2=\frac{-2-\sqrt{185}}{10} \]

OpenStudy (helder_edwin):

did u get all this?

OpenStudy (rayford):

yes is that it?

OpenStudy (helder_edwin):

yeah u can use a calculator to compute \(\sqrt{185}\) if u need it but that is it

OpenStudy (rayford):

thak you very much i understand this now :[)!

OpenStudy (helder_edwin):

great u r w